# Contrastive Self-Supervised Learning As Neural Manifold Packing

### v1

@@ 4.1 Linear evaluation
Following standard linear evaluation protocols, we froze the pretrained ResNet-50 backbone encoder and trained a linear classifier on top of the representation. Training was conducted for 100 epochs on ImageNet-1K and 200 epochs on ImageNet-100. Classification accuracies are reported on the corresponding validation sets (Table 1). We find that CLAMP achieves competitive performance relative to existing SSL algorithms and even sets a new state of the art on ImageNet-100. While performance can be further improved using 8 augmented views, as shown in the Appendix K. we limit our experiments to 4 views due to computational resource constraints. See Appendix E for details.
@@ 5 Training dynamics reflect geometrical changes in the embedding space
We analyze the dynamics of the manifold packing by tracking the evolution of the number of neighbors and manifold size during training on ImageNet-1K. Two sub-manifolds are counted as neighbors if the Euclidean distance between their centroids is smaller than the sum of their radii (i.e., $||q_{i}-q_{j}||_{2}<r_{i}+r_{j}$). For each epoch, we compute the average number of neighbors per sub-manifold (with $m=4$ views) and the average manifold size, quantified by $\mathbb{E}_{i\sim\mathrm{data}}[\sqrt{\Lambda_{i}}/m]$. Both metrics were evaluated at the end of each training epoch on the validation data, which comprises a random 1% split of the training data. As shown in Fig. 3(a) and (b) both metrics decrease during training. Fig. 3(c) illustrates that the embeddings start in a collapsed state due to random initialization but evolve into more structured representations with increased pairwise distances. This behavior is reminiscent of random organization models discussed in Sec. 2.1, where local particle density and inter-particle spacing exhibit similar temporal patterns [17, 15]. These dynamics reflect the growing similarity of positive samples and the increasing separation of negative samples, indicating improved feature discrimination over time.
Figure 3: Training dynamics: (a) Number of neighbours as the function of epochs. (b) Average embedding sub-manifold sizes as the function of epochs. (c) Distances between pair of embeddings for untrained and trained networks.
@@ 6 Properties of the representations
To quantify the geometric properties of the learned representations, we analyzed the CLAMP representation obtained for CIFAR-10 using a ResNet-18 backbone architecture. As shown in Appendix F, we achieve $\sim 90\%$ top-1 accuracy (without hyperparameter tuning) which is competitive with alternative SSL frameworks. For each sub-manifold, we computed its centroid and alignment vector (defined as the principal eigenvector of its covariance matrix), grouping them according to their class labels. We compared the geometric properties such as centroid distances, centroid cosine similarity, and alignment cosine similarity across inter-class and intra-class sub-manifold pairs. To ensure robust estimation, we randomly sampled 800 images from the test dataset and generated 100 augmentations per image to compute the manifold properties. This process was repeated 10 times.
@@
To demonstrate the emergence of category-specific manifolds, we visualize the representation space of the CIFAR-10 test dataset (without augmentation) using t-SNE  [44], a method that preserves local neighborhood structure while projecting high-dimensional data into a lower-dimensional space. As shown in Fig. 5, embeddings from different classes form distinct, well-separated clusters, indicating the formation of class-specific manifolds in the learned representation space. This clear clustering underscores the structured organization induced by CLAMP and demonstrates its effectiveness in producing representation that are separable in representation space.
Methods V1 V2 V4 IT
@@
MMCR (8 views) 0.270 0.311 0.577 0.554
  CLAMP (4 views) 0.258 ±$\pm 0.013$ 0.336 ±$\pm 0.017$ 0.558 ±$\pm 0.004$ 0.570 ±$\pm 0.005$
Table 2: We use public data provided by Brain-Score library from [45] for V1 and V2, from [46] for V4 and IT. The benchmarks for other methods are from [32].
@@ 7 Biological implications
Neuroscience has long served as a source of inspiration for advancements in artificial neural networks, prompting the question of whether SSL models can replicate response statistics observed in cortex. Stringer et al. [47] analyzed stimulus-evoked activity in mouse V1 (primary visual cortex) and found that the eigenspectrum of its covariance matrix follows a power-law decay, $\lambda\propto n^{-\alpha}$ with $\alpha\approx 1.04$, and further proved that $\alpha>1$ is indeeed necessary to ensure differentiability of the neural code. After training CLAMP on ImageNet-1K, we measured the eigenspectrum of its feature covariance and observed a power law decay $\lambda\propto n^{-1.013}$, closely matching the cortical exponent and theoretical predictions [47].
@@ 8 Discussion
We have introduced CLAMP, a novel self-supervised learning framework that recasts contrastive learning as a neural-manifold packing problem, guided by a physics-inspired loss. By promoting similarity of positive pairs and repulsion of negatives through purely local multi-view repulsive interactions, CLAMP achieves competitive, and in some cases state-of-the-art, performance across standard downstream benchmarks. Importantly, our analysis of training dynamics reveals a progressive structuring of the embedding space. Sub-manifolds that are initially overlapping evolve into well-separated compact structures. We further show that class-level manifolds emerge from the aggregation of sub-manifolds, which interact purely through repulsion, a phenomenon reminiscent of phase separation in statistical physics. The mechanism underlying this emergent organization should be elucidated in future theoretical work.
We demonstrate that CLAMP offers a compelling account of how complex visual features might self-organize in the cortex without explicit supervision, based on direct comparisons between its learned representations and experimental neural data. However, these findings should be interpreted with caution. Brain-Score’s linear mapping between deep nets output and experimental activity may overlook non-linear relationships between model and neural representations. Moreover, CLAMP’s use of large-batch contrastive optimization with backpropagation is unlikely to reflect biologically realistic learning mechanisms. Developing a local, Hebbian-style learning rule inspired by manifold packing is a critical direction for future work.
Our loss function leverages the sizes and locations of sub-manifolds within the embedding space. Incorporating orientation alignment among sub-manifolds may enable denser packing, potentially improving downstream task performance. However, we currently omit orientation information due to its high computational cost: accurately estimating a sub-manifold’s orientation requires on the order of $O(D)\sim 100$ points, $\sim 100$ augmentations per input, which is infeasible at the scale of CIFAR-10 or ImageNet.
@@ Appendix A Algorithm
         $y_{i}^{1},...y_{i}^{m}=f_{\theta}(x_{i}^{1}),...f_{\theta}(x_{i}^{m})$ extract features
         $z_{i}^{1},...z_{i}^{m}=g_{\phi}(y_{i}^{1}),...g_{\phi}(y_{i}^{m})$ projection
@@
     $\tilde{z}_{i}^{k}=(z_{i}^{k}-c)/\lVert z_{i}^{k}-c\rVert_{2}$ for $i=1...b,k=1...m$
@@
         $\Lambda_{i}=\frac{1}{m}\sum_{k}(\tilde{z}_{i}^{k}-q_{i})^{T}(\tilde{z}_{i}^{k}-q_{i})$ covariance matrix
         $r_{i}=r_{s}\sqrt{\frac{1}{m}\mathrm{Tr}(\Lambda_{i})}$ manifold size
@@
     $\mathcal{L}_{overlap}=\begin{cases}\sum\limits_{i\neq j}\left(1-\frac{\lVert q_{i}-q_{j}\rVert_{2}}{r_{i}+r_{j}}\right)^{2}\text{,if }\lVert q_{i}-q_{j}\rVert_{2}<r_{i}+r_{j}\\ 0\text{,\quad otherwise }\end{cases}$
@@ Appendix C Time complexity for the loss function
Note that only the diagonal part of the covariance matrix $\Lambda_{i}$ is required in the loss function, it is not necessary to calculate the complete covariance matrix. Therefore, each iteration, the time complexity for computing sub-manifold sizes $r_{i}$ is $O(bmD)$, the time complexity for the pairwise loss function is $O(b^{2}D)$, therefore the time complexity for each iteration is $O(bD\ \text{min}(m,b))$. Since the number of iteration is $N/b$ where $N$ is the size of the dataset, the complexity for calculating the loss function for one epoch is $O(ND\ \text{min}(m,b))$.
@@ Appendix D Self-supervised pretraining
We employ the LARS optimizer for self-supervised learning with its default trust coefficient (0.001). Our learning-rate schedule begins with a 10-step linear warm-up, after which we apply cosine decay. To enable fully parallel training, we convert all BatchNorm layers to synchronized BatchNorm and aggregate embeddings from all GPUs when computing the loss. In addition, we exclude every bias term and all BatchNorm parameters from weight decay and use weight decay parameter = $10^{-6}$. For CIFAR-10 dataset, we removed the max pooling layer and modified the first convolution layer to be kernel size=1,stride=1,padding=2. We use $m=16$ views for CIFAR-10 dataset and $m=4$ for ImageNet-1K/100.$r_{s}=7.0$ is applied for CIAR-10 and ImageNet-100/1K. Other parameters are shown in Table 5 where the training learning rate = lr $\times$ batch size/256.
@@
ImageNet-100 ResNet-50 [8192,8192,512] 800 2.0 0.9 $200$
ImageNet-1K ResNet-50 [8192,8192,512] 1024 0.9 0.9 $100$
@@ Appendix E Linear evaluation
CIFAR10 dataset: We use the Adam optimizer with learning rate =0.05×=0.05$\times$= 0.05 × and batch size/ 256.0 (batch size = 1024) to train the liner classifier for 100 epochs. We apply a cosine learning rate schedule.
ImageNet-100 dataset: We use SGD optimizer with Nesterov momentum and weight decay to train the linear classifier where the learning rate =0.05×=0.05$\times$= 0.05 × batch size / 256, batch size = 1024, momentum = 0.9 and weight decay = 1e-5. During training, we transform an input image by random cropping, resizing it to $224\times 224$, and flipping the image horizontally with probability 0.5. At test time we resize the image to $256\times 256$ and center-crop it to a size of $224\times 224$. We apply a cosine learning rate schedule to train the linear classifier for 100 epochs.
ImageNet-1K dataset:we use the same setup as ImageNet-100 but learning rate =1.6×=1.6$\times$= 1.6 × batch size / 256, batch size = 1024.
@@ Appendix F Linear evaluation on CIFAR10
Table 6: Linear evaluation on CIFAR10 dataset using resnet-18 as the backbone encoder. Results for SimCLR[1], SwAV [27], DeepCluster V2[27],Barlow twins[40] and BYOL [2] methods are take from solo-learn benchmarking [42].
@@ Appendix G Semi-supervised learning
We use semi-supervised learning via fine tuning on the backbone network with a linear classifier for 1% and 10% splits of the ImageNet-1K dataset. The same splits are applied as [1]. We use the SGD optimizer with momentum of 0.9 but no weight decay for cross entropy loss. We adopt the same data augmentation procedure during training and testing as in the linear evaluation protocol. We scale down the learning rate of the backbone parameters by a factor of 20. Both the backbone and classifier learning rates follow a cosine decay schedule over 20 epochs. We use a batch size of 256 and sweep the initial learning rate over {0.05, 0.1, 0.5}.
Appendix H Power law exponent for the eigen-spectrum
@@
Appendix I Brain-Score
@@
Appendix J Qualitative understanding of sub-manifolds in the embedding space
J.1 Manifold radius for singular $\Lambda_{i}$
@@
Even if $\Lambda_{i}$ is not a full-rank matrix, $\mathrm{Tr}(\Lambda_{i})=\lambda_{i}^{(1)}+...+\lambda_{i}^{(1)}$ is well defined. We get
@@
J.2 The relation between estimated sub-manifold radius and the estimated sub-manifold volume
@@
$\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}\left(\frac{1}{K}\sum_{i}{r_{s}\sqrt{\lambda_{i}^{(j)}}}\right)^{K}\leq\left(r_{s}\sqrt{\frac{\mathrm{Tr}(\Lambda_{i})}{K}}\right)^{K}\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}=\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}r_{i}^{K}$
@@
Appendix K Hyperparameters
We conducted an ablation study to study the role of different hyperparameters. Due to limited computational resources, $30\%$ of the training data from ImageNet-1K were selected and pretrained for 100 epochs (which is consistent with our baseline where the whole training set are pretrained for 100 epochs). We used the frozen-backbone linear evaluation accuracy as the criterion for ablation. We used lr = 8.0 $\times$ batch_size/256, weight decay parameter = 1e-6 and momentum = 0.0 for pretraining.
K.1 Scale parameter $r_{s}$
@@
Figure 7: Linear evaluation accuracy as the function of the size scale factor $r_{s}$. we use 4 views.
K.2 Number of augmentations $m$

### v2

@@ 4.1 Linear evaluation
Following standard linear evaluation protocols, we froze the pretrained ResNet-50 backbone encoder and trained a linear classifier on top of the representation. Training was conducted for 100 epochs on ImageNet-1K and 200 epochs on ImageNet-100. Classification accuracies are reported on the corresponding validation sets (Table 1). We find that CLAMP achieves competitive performance relative to existing SSL algorithms and even sets a new state of the art on ImageNet-100. Our model performance under linear evaluation depends only weakly on batch size but improves with an increased number of augmented views as shown in Appendix L. See Appendix E for details about linear evaluation.
@@ 4.2 Semi-supervised learning
4.3 Transfer to object detection tasks
To evaluate CLAMP’s generalizability beyond linear classification, we apply the pretrained model to object detection. Specifically, we report results using a Faster R-CNN architecture with a C4 backbone, pretrained with CLAMP on ImageNet-1K using 8 views and a batch size of 512 for 100 epochs. The model is fine-tuned on the VOC2007+2012 training set and evaluated on the VOC2007 test set (Table. 2). CLAMP delivers stronger detection performance than other self-supervised learning baselines, demonstrating that its learned manifold structure transfers effectively to spatially structured vision tasks. See Appendix H for details.
Methods mAP $\uparrow$ AP50 $\uparrow$ AP75 $\uparrow$
  SimCLR 54.4 81.6 61.0
Barlow Twins 53.1 80.9 57.7
BYOL 55.6 82.3 62.0
MoCo v2 54.7 81.7 60.2
MMCR 54.6 81.9 60.0
CLAMP 55.7 82.3 62.4
Table 2: Object detection on VOC datasets: The benchmarks (mAP, AP50 and AP75) for the baseline methods are from [34]. Models are pretrained on ImageNet-1K for 100 epochs before fine-tuning.
@@ 5 Training dynamics reflect geometrical changes in the embedding space
We analyze manifold packing dynamics by tracking the evolution of neighbor count and manifold size during ImageNet-1K training. Two sub-manifolds are considered neighbors if the Euclidean distance between their centroids is smaller than the sum of their radii ($q_{i}-q_{j}||2<r_{i}+r_{j}$). At each epoch, we compute the average number of neighbors (for $m=4$ views) and the average manifold size, $\mathbb{E}_{i\sim\mathrm{data}}[\sqrt{\Lambda_{i}}/m]$, using a 1% validation split. As shown in Fig. 3, both metrics decrease over training, indicating that embeddings evolve from an initial collapsed state toward more structured representations with increased pairwise distances. This mirrors random organization models [17, 15], where local density and spacing show similar temporal behavior, reflecting improved feature discrimination over time.
Figure 3: Training dynamics: (a) Number of neighbours as a function of epochs. (b) Average embedding sub-manifold sizes as the function of epochs. (c) Distances between pair of embeddings for untrained and trained networks.
@@ 6 Properties of the representations
To quantify the geometric properties of the learned representations, we analyzed the CLAMP representation obtained for CIFAR-10 using a ResNet-18 backbone architecture. As shown in Appendix F, we achieve $\sim 90\%$ top-1 accuracy (without hyperparameter tuning) which is competitive with alternative SSL frameworks. For each sub-manifold, we computed its centroid and alignment vector (defined as the principal eigenvector of its covariance matrix), grouping them according to their class labels. We compared the geometric properties such as centroid distances, centroid cosine similarity, and alignment cosine similarity of representations across inter-class and intra-class sub-manifold pairs. To ensure robust estimation, we randomly sampled 800 images from the test dataset and generated 100 augmentations per image to compute the manifold properties. This process was repeated 10 times.
@@
To demonstrate the emergence of category-specific manifolds, we visualize the representation space of the CIFAR-10 test dataset (without augmentation) using t-SNE  [46], a method that preserves local neighborhood structure while projecting high-dimensional data into a lower-dimensional space. As shown in Fig. 5, representations from different classes form distinct, well-separated clusters, indicating the formation of class-specific manifolds in the learned representation space. This clear clustering underscores the structured organization induced by CLAMP and demonstrates its effectiveness in producing representation that are separable in representation space.
Methods V1 $\uparrow$ V2 $\uparrow$ V4 $\uparrow$ IT $\uparrow$
@@
MMCR (8 views+ME) 0.270 0.311 0.577 0.554
  CLAMP (4 views) 0.258 $\pm 0.013$ 0.336 $\pm 0.017$ 0.558 $\pm 0.004$ 0.570 $\pm 0.005$
Table 3: Brain-score:We use public data from [47] for V1 and V2, and from [48] for V4 and IT to evaluate the brain-score [49, 50]. The benchmarks for all baseline methods are from [34]. In the MMCR method, ME refers to the momentum encoder.
@@ 7 Biological implications
Neuroscience has long served as a source of inspiration for advancements in artificial neural networks, prompting the question of whether SSL models can replicate response statistics observed in cortex. Stringer et al. [51] analyzed stimulus-evoked activity in mouse V1 (primary visual cortex) and found that the eigenspectrum of its covariance matrix follows a power-law decay, $\lambda\propto n^{-\alpha}$ with $\alpha\approx 1.04$, and further proved that $\alpha>1$ is indeed necessary to ensure differentiability of the neural code. After training CLAMP on ImageNet-1K, we measured the eigenspectrum of its feature covariance and observed a power law decay $\lambda\propto n^{-1.013}$, matching the cortical exponent and theoretical predictions [51].
@@ 8 Discussion
We have introduced CLAMP, a novel self‑supervised learning framework that recasts contrastive learning as a neural‑manifold packing problem, guided by a physics‑inspired loss. By promoting similarity of positive pairs and repulsion of negatives through purely short-range, multi-view repulsive interactions, CLAMP achieves competitive, and in some cases state-of-the-art, performance across standard downstream benchmarks. Importantly, our analysis of training dynamics reveals a progressive structuring of the embedding space. Sub‑manifolds that are initially overlapping evolve into well-separated compact structures. We further show that class-level manifolds emerge from the aggregation of sub-manifolds, which interact purely through repulsion, a phenomenon reminiscent of phase separation in statistical physics. The mechanism underlying this emergent organization should be elucidated in future theoretical work. It is also interesting to test if hierarchical cluster emerges during the learning process in the future.
We demonstrate that CLAMP offers a compelling account of how complex visual features might self-organize in the cortex without explicit supervision, based on direct comparisons between its learned representations and experimental neural data. However, these findings should be interpreted with caution. Brain-Score’s linear mapping between deep nets output and experimental activity may overlook non-linear relationships between model and neural representation. It also does not guarantee that CLAMP implements the same computations or dynamics as cortex. Moreover, CLAMP’s use of large-batch contrastive optimization with backpropagation is unlikely to reflect biologically realistic learning mechanisms. In contrast, Hebbian-style self-supervised learning frameworks with local update rules are considered biologically plausible [52]. Developing local self-supervised learning dynamics inspired by manifold packing, without relying on backpropagation, is an important direction for future research.
Our loss function leverages the sizes and locations of sub‑manifolds within the embedding space. Incorporating orientation alignment among sub‑manifolds may enable denser packing, potentially improving downstream task performance. However, we currently omit orientation information due to its high computational cost: accurately estimating a sub‑manifold’s orientation requires on the order of $O(D)\sim 100$ points, $\sim 100$ augmentations per input, which is infeasible at the scale of CIFAR‑10 or ImageNet.
@@ Appendix A Algorithm
   $y_{i}^{1},...y_{i}^{m}=f_{\theta}(x_{i}^{1}),...f_{\theta}(x_{i}^{m})$ # extract features
   $z_{i}^{1},...z_{i}^{m}=g_{\phi}(y_{i}^{1}),...g_{\phi}(y_{i}^{m})$ # projection
@@
  $\tilde{z}_{i}^{k}=(z_{i}^{k}-c)/\lVert z_{i}^{k}-c\rVert_{2}$ for i=1b,k=1mi=1...b,k=1...m
@@
   $\Lambda_{i}=\frac{1}{m}\sum_{k}(\tilde{z}_{i}^{k}-q_{i})^{T}(\tilde{z}_{i}^{k}-q_{i})$ # covariance matrix
   $r_{i}=r_{s}\sqrt{\frac{1}{m}\mathrm{Tr}(\Lambda_{i})}$ # manifold size
   # Note: in practice, only the diagonal elements of $\Lambda_{i}$ are calculated for $\mathrm{Tr}(\Lambda_{i})$.
@@
[…]
@@ Appendix C Time complexity for the loss function
Note that only the diagonal part of the covariance matrix $\Lambda_{i}$ is required in the loss function, it is not necessary to calculate the complete covariance matrix. Therefore, each iteration, the time complexity for computing sub-manifold sizes $r_{i}$ is $O(bmD)$, the time complexity for the pairwise loss function is $O(b^{2}D)$, therefore the time complexity for each iteration is O(bDmax(m,b))O(bD\ $\text${max}(m,b)). Since the number of iteration is $N/b$ where $N$ is the size of the dataset, the complexity for calculating the loss function for one epoch is O(NDmax(m,b))O(ND\ $\text${max}(m,b)).
@@ Appendix D Self-supervised pretraining
We employ the LARS optimizer for self‑supervised learning with its default trust coefficient (0.001). Our learning‑rate schedule begins with a 10‑step linear warm‑up, after which we apply cosine decay. To enable fully parallel training, we convert all BatchNorm layers to synchronized BatchNorm and aggregate embeddings from all GPUs when computing the loss. In addition, we exclude every bias term and all BatchNorm parameters from weight decay and use weight decay parameter = $10^{-6}$. For CIFAR-10 dataset, we removed the max pooling layer and modified the first convolution layer to be kernel size=1,stride=1,padding=2. We use $m=16$ views for CIFAR-10 dataset and $m=4,8$ for ImageNet-1K/100. $r_{s}=8.5$ is applied for CIFAR-10 and ImageNet-100/1K. Other parameters are shown in Table 6 where the base learning rate = lr $\times$ batch size/256.
@@
ImageNet-100 ResNet-50 [8192,8192,512] 512 2.0 0.9 $200$
ImageNet-1K ResNet-50 [8192,8192,512] 512 1.1 0.9 $100$
@@ Appendix E Linear evaluation
CIFAR10 dataset: We use the Adam optimizer with learning rate $0.05\times$ and batch size/ 256.0 (batch size = 1024) to train the linear classifier for 100 epochs. We apply a cosine learning rate schedule.
ImageNet-100 dataset: We use SGD optimizer with Nesterov momentum and weight decay to train the linear classifier where the learning rate $0.05\times$ batch size / 256, batch size = 1024, momentum = 0.9 and weight decay = 1e-5. During training, we transform an input image by random cropping, resizing it to $224\times 224$, and flipping the image horizontally with probability 0.5. At test time we resize the image to $256\times 256$ and center-crop it to a size of $224\times 224$. We apply a cosine learning rate schedule to train the linear classifier for 100 epochs.
ImageNet-1K dataset:we use the same setup as ImageNet-100 but base learning rate $1.6\times$ batch size / 256, batch size = 1024.
@@ Appendix F Linear evaluation on CIFAR10
Table 7: Linear evaluation on CIFAR10 dataset using resnet-18 as the backbone encoder. Results for SimCLR [1], SwAV  [29], DeepCluster V2 [29],Barlow twins [42] and BYOL [2] methods are taken from solo-learn benchmarking [44].
@@ Appendix G Semi-supervised learning
We use semi-supervised learning via fine tuning on the backbone network with a linear classifier for 1% and 10% splits of the ImageNet-1K dataset. The same splits are applied as [1]. We use the SGD optimizer with momentum of 0.9 but no weight decay for cross entropy loss. We adopt the same data augmentation procedure during training and testing as in the linear evaluation protocol. We scale down the learning rate of the backbone parameters by a factor of 20. Both the backbone and classifier use the cosine-decay learning rate schedule. We use a batch size of 256 and sweep the initial learning rate over {0.1, 0.2, 0.4, 0.6}.
Appendix H Object detection
we apply the Faster R-CNN detector with a R50-C4 backbone (pretrained on ImageNet-1K with 8 views and batch size 512 for 100 epochs), fine-tuned on VOC2007+2012 training dataset and tested on VOC2007 test dataset. Our implementation follows the procedure of [3], except that we set the base learning rate to 0.07.
Appendix I Power law exponent for the eigen-spectrum
@@
Appendix J Brain-Score
@@
Appendix K Qualitative understanding of sub-manifolds in the embedding space
K.1 Manifold radius for singular $\Lambda_{i}$
@@
Even if $\Lambda_{i}$ is not a full-rank matrix, $\mathrm{Tr}(\Lambda_{i})=\lambda_{i}^{(1)}+...+\lambda_{i}^{(K)}$ is well defined. We get
@@
K.2 The relation between estimated sub-manifold radius and the estimated sub-manifold volume
@@
$V=\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}\left(\frac{1}{K}\sum_{i}{r_{s}\sqrt{\lambda_{i}^{(j)}}}\right)^{K}\leq\left(r_{s}\sqrt{\frac{\mathrm{Tr}(\Lambda_{i})}{K}}\right)^{K}\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}=\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}r_{i}^{K}$
@@
Appendix L Hyperparameters
We conducted an ablation study to investigate the role of different hyperparameters. Due to limited computational resources, ablation study on parameter $r_{s}$ and number of views $m$, $30\%$ of the training data from ImageNet-1K were selected and pretrained for 100 epochs (which is consistent with our baseline where the whole training set are pretrained for 100 epochs). For the ablation on batch size, we employed the entire ImageNet-1K dataset. We used the frozen-backbone linear evaluation accuracy on the entire test dataset as the criterion for ablation.
L.1 Scale parameter $r_{s}$
@@
Figure 7: Linear evaluation accuracy as the function of the size scale factor $r_{s}$ on $30\%$ ImageNet-1K dataset. we use 4 views.
L.2 Number of augmentations $m$